OpenStudy (phebe):

Phythagreon Therom help!! XD @Sparrowhawk69

OpenStudy (anonymous):

I can help!

OpenStudy (phebe):

ok thanx @NERD85176

OpenStudy (anonymous):

I'm pretty good at that. What's the question?

OpenStudy (phebe):

lol wat?? @natedog12

OpenStudy (the_fizicx99):

Um, does that mean A = 100 B = ? C = 125 ? Find B?

OpenStudy (phebe):

its a right angle

OpenStudy (phebe):

yes @tHe_FiZiCx99

OpenStudy (paki):

just use Pythagoras theorem and solve...

OpenStudy (anonymous):

Were is the question???

OpenStudy (anonymous):

Oh nvm lol

OpenStudy (anonymous):

\[a ^{2}+ b ^{2} = c ^{2}\] Is the formula, so first, plug in what you know.

OpenStudy (the_fizicx99):

Yah. The formula is: a^2 + b^2 = c^2 Plug in your values, 100^2 + b^2 = 125^ 10,000 + b^2 = 15,625 Subtract 10,000 from both sides, you'll be left with b^2 = ? Then take the square root of both sides and that's b.

OpenStudy (anonymous):

@natedog12 Ask that in the chat, don't go posting irrelevant stuff in people's questions!

OpenStudy (anonymous):

The answer is obviously 25

OpenStudy (phebe):

u sure @Sparrowhawk69

OpenStudy (the_fizicx99):

o_o Not even close. ^^

OpenStudy (anonymous):

No it is not 25.

OpenStudy (phebe):

so that answer is wrong @tHe_FiZiCx99

OpenStudy (the_fizicx99):

Yes it was wrong... Can you solve it?

OpenStudy (phebe):

um... i guess so so wat u tthink the answer is

OpenStudy (anonymous):

We know the answer but we want you to solve it. :)

OpenStudy (the_fizicx99):

Doesn't matter what I think the answer is, what do YOU think the answer is.

OpenStudy (phebe):

okkkkkay

OpenStudy (paki):

the answer is "75"...

OpenStudy (anonymous):

Just subtract the a value from the c value.

OpenStudy (paki):

pleasure...

OpenStudy (anonymous):

@paki Phebe needed to do it so she could understand it.

OpenStudy (the_fizicx99):

Smh.

OpenStudy (anonymous):

^

OpenStudy (paki):

@mipie13 bro you guys already did all... :)

OpenStudy (anonymous):

O_o.. We didn't do all of it.

OpenStudy (anonymous):

We were taking her through it step by step so she could understand.

OpenStudy (phebe):

i already do understand i jus wanted the answer for that one

OpenStudy (the_fizicx99):

If you understood it why didn't you answer it yourself?

OpenStudy (phebe):

becuz i was in a hurry

OpenStudy (the_fizicx99):

It took longer for me to explain, you could've done this easily with a calculator ._.

OpenStudy (phebe):

ik and i got a calculater thank all yall for the help okayyy

OpenStudy (anonymous):

o_e...

OpenStudy (anonymous):

It took me like a minute to do this with a calculator..